[Hallc_running] [New Comment] RC Daily Update

rmontgom at jlab.org rmontgom at jlab.org
Mon Nov 26 11:50:07 EST 2018

 - by rmontgom on Mon, 11/26/2018 - 11:47.

RC daily update addition - notice for workers in Hall - NB: If you are working in the Hall today please be aware that maintenance are going to install a lift close to the Hall entrance in attempt to fix the flooding.


Logentry Text:
- Smooth beam since yesterday (https://logbooks.jlab.org/entry/3631741).
- Day yesterday completed Group 18 (Q2 = 5.5, x = 0.5, pi+,  pi-).
- During configuration change to Group 3 (Q2 = 4, x = 0.4, pi+) HMS dipole tripped resulting in ~2hr related downtime (DTM 5528). After interlock reset, it could not be ramped to MOL without tripping. Dipole was ramped by hand to target value, field settled ok. The issue needs to be fixed - S. Lassiter is investigating.
- Swing completed Group 3 and moved to Group 4 (Q2 = 4, x = 0.4, pi-).
- Owl completed Group 4 and moved to collect data for SIDIS pt experiment E12-09-17 (https://hallcweb.jlab.org/wiki/index.php/5_pass_Run_Plan#Kinematic_Group_15) at settings 16-11 and 15-11.
- Beam was shut off 8am. Rad Con sweeped the Hall which is now in restricted access.

RC Meeting:
- Since we have no beam we can cancel the RC run plan meeting today (unless there is a special request, then I will certainly schedule). The summary of running since last week will be discussed at today's Hall C meeting. We will have an RC meeting tomorrow to plan the start of the upcoming Kaon LT experiment.

Notices in effect for shift crews:
* Shift changes. TO needed for all shifts. SL cancelled Mon day 11/26, Mon swing 11/26, Tue owl 11/27, Tue day 11/27. SL on standby Tue swing 11/27, Wed owl 11/28 and Wed day 11/28. SL shifts normal from Wed swing 11/28 currently. If on standby please be ready to come to counting house. We will know more Tue/Wed about physics beam delivery - stay tuned for updates.

Planned accelerator configuration changes:
Monday, Nov 26: Currently down for energy change to 1.86 GeV/pass. We will be moved to 500MHz laser. MCC want Hall locked up start of Tue swing to start beam studies.

Planned accesses to the Hall:
Monday, Nov 26: Work in Hall should be coordinated with W. Kellner. Aerogel change by V. Berdnikov and Walter's team on-going. HGC work by B. Sawatzky and University of Regina team ongoing. D. Meekins target loop 3 JT valve issue on-going. V. Kumar (Regina group) walkthrough today needs time TBC. S. Lassiter is looking into HMS dipole issue/rotation issues.
Tuesday, Nov 27: possible installation of brushes near spectrometer wheels for S. Lassiter/W. Kellner.
Please let me know any other plans for work in the Hall Mon/Tue.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3631779
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181126/318b4740/attachment-0001.html>

More information about the Hallc_running mailing list